# Dynamic angular synchronization under smoothness constraints

### v2

@@ Appendix B Proof of Lemma 3
which we will bound via the Hanson-Wright inequality [39, Thm. 1.1]. To this end, we first derive an upper bound on $\mathbb{E}[u^{\top}(I_{n}\otimes\mathcal{P}_{\tau,n})u]$ as follows. Denoting $Z_{1,\operatorname{R}}(k)$ (resp. $\overline{G}_{1,\operatorname{R}}^{*}(k)$) to be the real part of $Z_{1}(k)$ (resp. $\overline{G}_{1}^{*}(k)$), we obtain
@@ E.1 RMSE versus noise-level
Figure 7: RMSE versus $\gamma(=\sigma)$ for the AGN model ($n=30$, $T=100$) with $S_{T}\in\left\{{1/T,1,T^{1/4}}\right\}$, results averaged over $20$ MC runs. We choose $\lambda_{\text{scale}}=10$ for GTRS-DynSync. Results on left show average RMSE, the right panel shows average RMSE $\pm$ standard deviation.
@@
Figure 8: RMSE versus $\gamma(=\eta)$ for the Outliers model ($n=30$, $p=0.2$, $T=100$) with $S_{T}\in\left\{{1/T,1,T^{1/4}}\right\}$, results averaged over $20$ MC runs. We choose $\lambda_{\text{scale}}=10$ for GTRS-DynSync. Results on left show average RMSE, the right panel shows average RMSE $\pm$ standard deviation.
@@ E.2 Simulations for a Projected Power Method (PPM)
[…]
@@
Figure 9: RMSE versus $\gamma(=\sigma)$ for PPM-DynSync for the AGN model ($n=30$, $T=100$) with $S_{T}=1/T$ and different initializations. Results averaged over $20$ MC runs. We choose $\lambda_{\text{scale}}=10$ for GTRS-DynSync and PPM-DynSync.

### v3

@@ Appendix B Proof of Lemma 3
which we will bound via the Hanson-Wright inequality [40, Thm. 1.1]. To this end, we first derive an upper bound on $\mathbb{E}[u^{\top}(I_{n}\otimes\mathcal{P}_{\tau,n})u]$ as follows. Denoting $Z_{1,\operatorname{R}}(k)$ (resp. $\overline{G}_{1,\operatorname{R}}^{*}(k)$) to be the real part of $Z_{1}(k)$ (resp. $\overline{G}_{1}^{*}(k)$), we obtain
@@ E.1 RMSE versus noise-level
Figure 7: RMSE versus $\gamma(=\sigma)$ for the AGN model ($n=30$, $T=100$) with $S_{T}\in\left\{{1/T,1,T^{1/4}}\right\}$, results averaged over 2020 MC runs. We choose $\lambda_{\text{scale}}=10$ for GTRS-DynSync. Results on left show average RMSE, the right panel shows average RMSE $\pm$ standard deviation.
@@
Figure 8: RMSE versus $\gamma(=\eta)$ for the Outliers model ($n=30$, $p=0.2$, $T=100$) with $S_{T}\in\left\{{1/T,1,T^{1/4}}\right\}$, results averaged over 2020 MC runs. We choose $\lambda_{\text{scale}}=10$ for GTRS-DynSync. Results on left show average RMSE, the right panel shows average RMSE $\pm$ standard deviation.
@@ E.2 Simulations for a Projected Power Method (PPM)
For initialization, we will employ $\texttt{GTRS-DynSync},{\texttt{GMD-LTRS-DynSync}}$, LTRS-GS-DynSync and the naive spectral algorithm considered in the experiments (Section 4). For PPM-DynSync, the parameter $\lambda$ is chosen via the same grid search procedure as for GTRS-DynSync (see Section 4); we chose $\lambda_{scale}=10$ for both PPM-DynSync and GTRS-DynSync. Surprisingly, we found that PPM-DynSync does not improve the performance of the seed, and in general worsens the estimation error. While a more detailed investigation of PPM-DynSync is left for future work, we illustrate its performance in Fig. 9 for the AGN model. Similar behaviour was observed for the Outliers model, and this was the case for different choices of $\lambda_{scale}$ for PPM-DynSync.
@@
Figure 9: RMSE versus $\gamma(=\sigma)$ for PPM-DynSync for the AGN model ($n=30$, $T=100$) with $S_{T}=1/T$ and different initializations. Results averaged over 2020 MC runs. We choose $\lambda_{\text{scale}}=10$ for GTRS-DynSync and PPM-DynSync.